The side of a perfect cube of sugar is 0.75 inches. Answer the following questions:

A. Report the length of the cube in cm and mm (show your work)

B. What is the volume of the cube? (show your work)

C. What is the mass of the cube if the density is sucrose (table sugar) is 1.59g/cm3?

D. How many moles of sucrose (C12H22O11) are there in the cube.
Answer #1

A)

1 inch is 25.4 mm

we have:

L = 0.75 inch

= 0.75*25.4 mm

= 19.05 mm

= 1.905 cm

Answer: 19.05 mm and 1.905 cm

B)

Volume = L^3

= (1.905 cm)^3

= 6.91 cm^3

Answer: 6.91 cm^3

C)

use:

mass = density * volume

= 1.59 g/cm^3 * 6.91 cm^3

= 11.0 g

Answer: 11.0 g

D)

Molar mass of C12H22O11,

MM = 12*MM(C) + 22*MM(H) + 11*MM(O)

= 12*12.01 + 22*1.008 + 11*16.0

= 342.296 g/mol

mass(C12H22O11)= 11.0 g

use:

number of mol of C12H22O11,

n = mass of C12H22O11/molar mass of C12H22O11

=(11 g)/(3.423*10^2 g/mol)

= 3.214*10^-2 mol

Answer: 3.21*10^-2 mol